### v1

@@ 5.3 Physical Priors Injection
  • Injecting into the model (the circle labelled 2 in Fig. 2): Integrating Physical Priors into the model involves incorporating physical equations, such as partial differential equations (PDEs), directly into the model architecture. One effective approach is the use of Physics-Informed Neural Networks (PINNs), where the neural network is trained not only to minimize the data-driven loss but also to satisfy the governing PDEs that describe the underlying physical processes. This allows the model to leverage both data and physical knowledge, leading to more accurate and physically consistent predictions, especially in scenarios where data is sparse or noisy.
  • Injecting into the inputs (the circle labelled 3 in Fig. 2): Integrating physical priors $\mathcal{K}_{in}$ into the model’s raw inputs $X_{in}$ involves incorporating meteorological knowledge or historical data as additional input signals to guide the model. For example, we can introduce weather patterns, such as wind speed, humidity, or pressure distributions, derived from historical climate data or scientific literature, as supplementary features concatenated with the raw inputs $[X_{in},\mathcal{K}_{in}]$. These physical inputs provide the model with prior knowledge of atmospheric conditions, enabling it to make more informed predictions. By including these established physical signals, the model is better equipped to understand the underlying dynamics of climate systems, leading to improved accuracy and consistency in its outputs, particularly in forecasting and long-term climate projections.
@@ 5.4 Scalability in Large-Scale Systems
Large-scale and complex systems, such as global climate models (GCMs) [2], form the cornerstone of modern atmospheric research. These models integrate vast amounts of observational and simulated data to study Earth’s climate, making them critical tools for understanding phenomena ranging from regional weather patterns to global climate change. However, their computational demands, complex parameterizations, and inherent uncertainty present ongoing challenges.
@@
Another crucial aspect of scalability lies in the adaptability of the proposed model. As a pretrained foundation model, it can be fine-tuned for various tasks without requiring extensive retraining. For example, after training on a global dataset, the model can be adapted to regional scales or specific phenomena, such as monsoon dynamics or Arctic sea ice conditions. This ability to efficiently fine-tune on new datasets or emerging scientific challenges makes the model highly versatile, allowing researchers to address a wide range of atmospheric science problems with reduced computational effort and time.
@@
6 Conclusion and Future Directions
This work presents a novel approach to weather and climate sciences by leveraging a physics-guided multimodal transformer framework, which integrates diverse meteorological data modalities while incorporating critical physical principles. Through the analysis of key tasks, such as weather prediction, numerical super-resolution, and climate recognition, we show how transformer models can be guided by physical knowledge to achieve superior performance in atmospheric science problems. By representing different atmospheric layers (e.g., high-altitude and low-altitude data) as separate modalities, this approach tokenizes and processes these inputs in a manner that enhances model understanding. The transformer’s attention mechanism serves as an effective tool for fusing these diverse data streams, while the integration of physical knowledge further improves the predictive accuracy and scientific validity of the model.
Key insights and directions for future research include:
  1. 1.
    Enhancing Dataset Quality for Pretraining: The foundation of any AI-driven weather or climate model is the quality and alignment of its data. The development of high-quality, comprehensive datasets that represent the diverse modalities of climate data is crucial. These datasets should be designed with alignment methods that facilitate effective pretraining of transformer models. The inclusion of both observational data and simulated data, with careful consideration of physical consistency across modalities, will provide a solid basis for building stronger base models capable of better generalization.
  2. 2.
    Refining Multimodal Data Alignment and Unified Encoding: Weather and climate sciences often involve multiple sources of data, each with different resolutions, time scales, and modalities (e.g., satellite imagery, ground sensors, atmospheric models). Future research should focus on developing better alignment methods to bring these multimodal data sources into a unified encoding scheme. A well-designed encoding strategy will allow the model to capture cross-modal dependencies more effectively and leverage the strengths of each modality—such as spatial data from satellite imagery and temporal data from weather sensors—while ensuring consistency with underlying physical laws.
  3. 3.
    Addressing Challenges with Noisy and Incomplete Data: Meteorological data is frequently affected by noise, sensor errors, and missing values, which poses significant challenges for training robust AI models. Given the importance of maintaining physical consistency in weather and climate models, it is essential to explore how AI models can handle noisy labels and incomplete data in a way that preserves the underlying physical constraints. Techniques like semi-supervised learning, noise-robust training, and data imputation, combined with the physics-guided approach, can help mitigate these challenges and produce more reliable models in real-world climate science scenarios.
  4. 4.
    Improving Model Inference Speed through Optimization: The increasing complexity of AI models, especially in the context of weather and climate sciences, necessitates improvements in inference speed. Optimizing the model for faster response times without sacrificing performance is key for real-time applications such as climate monitoring, disaster prediction, and decision support. Model compression techniques, such as pruning, quantization, and knowledge distillation, could play an important role in this regard. Additionally, the physics-guided approach allows for selective simplifications of certain processes, which may enable more efficient computation without losing critical physical accuracy.
In conclusion, the physics-guided multimodal transformer framework represents a promising pathway for advancing weather and climate sciences through AI. By grounding the transformer model in physical principles and ensuring the integration of diverse climate data modalities, we can create more accurate, interpretable, and efficient models for addressing the challenges of weather and climate prediction along with monitoring. Moving forward, future research should focus on improving data alignment, enhancing the incorporation of physical knowledge, and addressing practical challenges such as data quality and model efficiency. The potential of AI in weather and climate sciences is vast, and by guiding these models with physics, we can ensure that they are not only powerful but also scientifically robust.

### v2

@@ 4.3 Physical Priors Injection
  • Injecting into the model (the circle labelled 2 in Fig. 2): Integrating Physical Priors into the model involves incorporating physical equations, such as PDEs, directly into the model architecture. One effective approach is the use of Physics-Informed Neural Networks (PINNs), where the neural network is trained not only to minimize the data-driven loss but also to satisfy the governing PDEs that describe the underlying physical processes. This allows the model to leverage both data and physical knowledge, leading to more accurate and physically consistent predictions, especially in scenarios where data is sparse or noisy.
  • Injecting into the inputs (the circle labelled 3 in Fig. 2): Integrating physical priors $\mathcal{K}_{in}$ into the model’s raw inputs $X_{in}$ involves incorporating meteorological knowledge or historical data as additional input signals to guide the model. For example, we can introduce weather patterns, such as wind speed, humidity, or pressure distributions, derived from historical climate data or scientific literature, as supplementary features concatenated with the raw inputs $X_{in},\mathcal{K}_{in}]$. These physical inputs provide the model with prior knowledge of atmospheric conditions, enabling it to make more informed predictions. By including these established physical signals, the model is better equipped to understand the underlying dynamics of climate systems, leading to improved accuracy and consistency in its outputs, particularly in forecasting and long-term climate projections.
@@ 4.4 Scalability in Large-Scale Systems
Large-scale and complex systems, such as global climate models (Stute et al., 2001), form the cornerstone of modern atmospheric research. These models integrate vast amounts of observational and simulated data to study Earth’s climate, making them critical tools for understanding phenomena ranging from regional weather patterns to global climate change. However, their computational demands, complex parameterizations, and inherent uncertainty present ongoing challenges.
@@
Another crucial aspect of scalability lies in the adaptability of the proposed path. As a pretrained foundation model, it can be fine-tuned for various tasks without requiring extensive retraining. For example, after training on a global dataset, the model can be adapted to regional scales or specific phenomena, such as monsoon dynamics or Arctic sea ice conditions. This ability to efficiently fine-tune on new datasets or emerging scientific challenges makes the model highly versatile, allowing researchers to address a wide range of atmospheric science problems with reduced computational effort and time.
@@
5 Alternative Views
While we advocate for the integration of physical laws into multimodal transformers, we acknowledge and address two prominent alternative perspectives in the current machine learning community.
5.1 The Symmetry of Scaling: Can Data Alone Supplant Explicit Physics?
A significant counter-argument, rooted in the Scaling Law hypothesis (Kaplan et al., 2020), suggests that sufficiently large models trained on massive datasets can implicitly “discover” physical principles without explicit guidance (Lam et al., 2023). Proponents of this view might argue that hard-coding physical constraints acts as a form of inductive bias that may ultimately limit the model’s capacity to represent complex, non-linear atmospheric dynamics that are not yet fully captured by classical equations.
Our Response: While scaling has led to impressive performance, “implicit discovery” of physics is often fragile. In scenarios involving out-of-distribution events, such as unprecedented extreme weather driven by climate change, purely data-driven models frequently produce physically impossible outputs (e.g., negative precipitation or violations of mass conservation). Explicit physics-guidance ensures that the model remains within the manifold of reality, providing a safety floor that scaling alone cannot guarantee when historical data is no longer a reliable predictor for future.
5.2 The Computational Trade-off: Accuracy vs. Physical Consistency
Another credible alternative view posits that the integration of physical constraints, especially through complex loss functions or modified attention kernels, introduces significant computational overhead and optimization challenges. Critics may argue that in operational forecasting, where inference speed is paramount, a faster, “approximately correct” data-driven model is more practically valuable than a slower, “physically perfect” transformer.
Our Response: We contend that the perceived trade-off between speed and consistency is a false dichotomy. By leveraging the multimodal transformer’s ability to fuse data efficiently, physical constraints can be used to reduce the search space during training, potentially leading to faster convergence. Furthermore, our proposed paradigm advocates for “physics-in-architecture” (e.g., tailored encoding) rather than merely adding expensive differential equation solvers to the inner loop of inference. The goal is to achieve “Physics-by-Design”, which maintains operational efficiency while ensuring scientific reliability.
6 Conclusion and Call to Action
This paper has argued that the current fragmented approach to meteorological AI is hitting a wall of scientific unreliability. We contend that the path forward lies in the adoption of a physics-guided multimodal transformer framework as the unified foundation for the field. By moving beyond “black-box” curve fitting and toward architectures that treat physical laws as structural imperatives, the machine learning community can provide the reliability and scientific consistency that critical climate policy and disaster response demand.
To realize this vision, we propose the following call to action for the meteorological AI community:
1) Develop Physics-Consistent Benchmarks: We call for a transition from standard error metrics (e.g., MSE) toward evaluation suites that explicitly penalize violations of physical conservation laws. High-quality, multimodal datasets must be co-designed by ML researchers and meteorologists to ensure that accuracy is always grounded in physical reality.
2) Prioritize “Physics-by-Design”: We advocate for a shift from treating physics as secondary input signals to embedding domain knowledge directly into transformer structures. Future research should prioritize architectures where physical constraints—such as fluid dynamics or thermodynamic limits—are integrated into attention mechanisms and encoding schemes, making consistency an inherent property of the model.
3) Ensure Operational Robustness and Efficiency: For these advanced frameworks to be viable, they must maintain physical plausibility even under data-sparse conditions (e.g., sensor failure). Furthermore, we urge the community to integrate model optimization techniques (such as quantization and distillation) to ensure that scientific rigor does not come at the cost of the real-time inference speed required for operational forecasting.
In conclusion, the success of AI in weather and climate sciences depends on its alignment with the immutable laws of physics. By guiding the next generation of multimodal transformers with these principles, we can ensure that AI serves as a scientifically robust pillar for understanding and protecting our planet’s future.
Appendix A Categories of AI Models in Weather and Climate Sciences
A.1 Forecasting Models
Forecasting models in weather and climate sciences predict future meteorological conditions using historical data. Convolutional Neural Networks (CNNs) and Graph Neural Networks (GNNs) have been widely used for their ability to capture spatial dependencies and model complex meteorological relationships. CNNs are effective in identifying spatial patterns in variables like temperature and pressure, while GNNs handle dynamic, grid-based weather system data across regions (Sønderby et al., 2020; Lam et al., 2023). For instance, MetNet combines CNNs, LSTMs, and autoencoders to capture multi-scale spatial and temporal features for short-term forecasts (Sønderby et al., 2020). Similarly, GraphCast employs GNNs within an “encode-process-decode” framework to model intricate spatiotemporal relationships, enhancing flexibility and accuracy over traditional approaches (Lam et al., 2023).
However, CNN and GNN-based models often fail to capture long-range temporal dependencies and global interactions crucial for accurate climate forecasting. Transformer-based models, leveraging attention mechanisms, address these limitations by focusing on relevant features across space and time (Bi et al., 2022; Pathak et al., 2022; Sun et al., 2024). Notable examples include PanGu Weather (Bi et al., 2022), which uses a 3D-Swin Transformer for spatiotemporal data processing, and FourCastNet (Pathak et al., 2022), integrating adaptive Fourier neural operators with vision Transformers for high-resolution data. These models better capture complex dependencies and model both local and global features effectively. These developments indicate a shift towards more scalable and flexible solutions, improving predictions of extreme weather events and long-term climate changes.
To optimize these models, we examine the classical objective function used in forecasting tasks. For example, the objective for PanGu Weather (Bi et al., 2022) is:
$\mathcal{L}=\text{MSE}(X_{t+1},f(X_{t},\theta)),$ (6)
$X_{t+1}\sim P(X_{t+1}|X_{1},\cdots,X_{t}),$ (7)
A.2 Classification and Detection Models
Classification and detection models in weather and climate sciences identify specific phenomena, such as extreme weather events, by classifying or localizing occurrences like storms and heatwaves. These models improve localization accuracy and address limited labeled data by extracting meaningful patterns from historical data for future predictions (Liu et al., 2016; Racah et al., 2017; Chen et al., 2019). Deep learning architectures, particularly CNNs, have been widely adopted for climate pattern detection and extreme event classification (Liu et al., 2016; Racah et al., 2017). Several works combining CNNs with recurrent structures like LSTMs capture the spatio-temporal dynamics of events such as typhoons (Chen et al., 2019).
$\mathcal{L}=-\sum_{i=1}^{N}\overline{y}_{t}\log(f(X_{t},\theta)),$ (8)
These tasks can be generalized as next-token prediction problems:
$Y_{t}\sim P(Y_{t}\mid X_{1},\cdots,X_{t}),$ (9)
A.3 Re-scaling Models
Re-scaling techniques, especially super-resolution (SR), enhance the resolution and quality of meteorological data by generating high-resolution outputs from lower-resolution inputs, crucial for accurate climate modeling (Sluiter, 2009; Vandal et al., 2017; Cheng et al., 2020; Watson et al., 2020). Deep learning models, particularly CNNs, effectively learn mappings between low and high-resolution data, capturing spatial and temporal patterns to improve dataset granularity. A primary application is downscaling precipitation data from coarse (e.g., 100 km) to fine resolutions (e.g., 12.5 km), better representing localized phenomena like extreme events and supporting more accurate weather predictions and climate assessments.
The objective function for re-scaling tasks typically uses MSE:
$\mathcal{L}=\text{MSE}(X_{\text{HR}},f(X_{\text{LR},\theta})),$ (10)
where $X_{\text{HR}_{t}}$ is the high-resolution data at time $t$.
More generally, re-scaling can be expressed as:
$Y_{t}\sim P(Y_{t}\mid X_{1},\cdots,X_{t}),$ (12)
A.4 Universal Embedding
Universal embedding techniques represent diverse meteorological data modalities within a common space, enabling integration across various meteorological-related tasks. This unified approach enhances model flexibility and adaptability (Kratzert et al., 2019; Man et al., 2023). Deep learning architectures, such as LSTMs and Masked AutoEncoders, have proven effective in creating universal embeddings. For example, Kratzert et al. (Kratzert et al., 2019) used EA-LSTM to model rainfall–runoff processes, while W-MAE (Man et al., 2023) employed self-supervised pre-training on ERA5 datasets, improving meteorological variable predictions and enhancing models like FourCastNet.
The objective for universal embedding is:
where $E_{t}$ is the embedding at time $t$, and $X,Y,Z,\cdots]_{1},\cdots,[X,Y,Z,\cdots]_{t}$ are inputs from various modalities. This framework integrates multiple data sources into a unified space, supporting diverse tasks like forecasting, classification, and anomaly detection.
A.5 Relationship Mining
Relationship mining in climate analysis explores connections between climatic factors, essential for understanding climate dynamics. Early methods used graph-based techniques to identify teleconnections (Kawale et al., 2013), while recent studies employ GNNs for more effective modeling (Romanova, 2023; Chishtie et al., 2024; Bhandari et al., 2024).
$G^{\prime}_{t}\sim P(G^{\prime}_{t}\mid G_{1},G_{2},\cdots,G_{t})$ (15)
A.6 Others